6/5/99 owl shift summary

Jason Hamm (jhamm@physics.arizona.edu)
Sat, 5 Jun 1999 05:06:12 -0700 (MST)

Shift Summary for owl shift, 06/05/99

On Shift: P. Shanahan and J. Hamm

Hours of beam delivered at all intensities: 0
Hours of special runs: ~4.5

Accounting of lost time: KTeV on access.

===================================================================
Runs ended this shift:

13366 : cosmic run
13367 - 13371 : laser scans
13372 - 13373 : cosmic runs
===================================================================
Accesses during shift: Open access for all shift.
===================================================================
Description of shift:

23:50 Cosmic run 13366 stopped to initialize DA for laser scan. Start of
run delayed due to problems with the installation of the new ET
interlock module.

2:00 New ET interlock installed and tested. Begin laser scan (13367)
to examine HCC illumination and search for hot HCC bits. Permit
lost in stream 5 repeatedly; laser crate repeater power cycled
remotely. This fixes the problem.

2:30 ET interlock trips. The LED on the interlock module indicates
the trip was caused by the temperature drifting outside the defined
limits. It is suspected that the high temperature threshold on the
new interlock module needs to be raised by a few degrees in order
match the threshold of the old module. Interlock is reset.

2:40 ET interlock trips again. Digital thermo-hydrometers report that
the maximum temperature seen by ET is 90 F. High temperature
threshold raised from 91 to 93 F to allow for fluctuations. No
more trips are seen during this shift. Remainder of run 13337 and
laser scan runs 13368, 13369 used by P. Shanahan for DPMT timing
studies.

4:00 Laser scan run 13370 started to study HCC bit illumination, search
for hot HCC bits, now that ET is stable. No missing HCC bits are
observed. Out of 7.2K laser triggers, no hot bits are seen.

4:30 P. Shanahan notices that the TRD_GAS.TCL window has been crashing
repeatedly. Problem fixed by rebooting "Serious Alarms" xterminal.

02:40- CsI timing work: runs 13367-13373
07:00 Needed to compensate effect of gain matching on signal timing.
Complete DBC scan --> new DBC skews loaded into pipeline. Timing
RMS goes from ~0.7ns to ~0.38ns. L1 delay adjusted again, globally
and for a few individual clock fanout channels, to maximize "bad
capid" free window.